if a household's income falls from R12 000 to R10 000, and its consumption falls from R9 500 to R8 000, then:
1
Expert's answer
2020-10-02T06:10:05-0400

MPC= Change in consumption/Change in income

Change in consumption= (9500-8000) 1500

Change in income = (12000-10000) 2000

MPC= 1500/2000= 0.75


Marginal propensity to consume is 0.75